### v1

@@ 2.1 Major Types of Awareness Emergence in Modern LLMs
SA represents the perception, comprehension, projection, and prediction of the future of the entities in the environment [25, 26, 27]. It was challenging to properly give a conclusive range of situational awareness, as many detailed manifestations differ across psychology, engineering, and systems ergonomics [28, 27, 29]. Situational awareness in LLMs typically refers to a model’s understanding of its identity and the context in which it is operating. It combines the monitoring of the environment and self-locating knowledge [8, 30, 31]. Primarily, machines and models possess knowledge about facts in the environment, their own influence in the world, and their understanding of how they will be influenced by the world. When utilizing their situational awareness, agents are perceiving the world state and their own state in it in real time. In AI safety literature, this concept is often defined as an LLM being aware that it is a model and recognizing whether it is currently in a testing scenario or deployed in the real world [8]. This form of awareness is critical because a highly capable model that knows it is being evaluated could potentially manipulate its behavior to appear aligned during tests, only to act differently once deployed. Early research has thus focused on measuring and inducing such awareness. Out-of-context reasoning evaluations show that large models can recall details about a test environment from training data (without explicit prompts) to succeed in that environment—an ability that strengthens with model size, hinting at emergent situational awareness as models scale [8]. Laine et al. [30, 31] performed a series of situational awareness tests and verified its importance in enhancing AI functions.
In this article, we distinguish situational awareness from self-awareness in that the content of SA is strictly restricted to sources of information that are not directly related to the agency. From a broader perspective, an agent will naturally know that it is in the scenario modeled by LLM, and this awareness involves both self-understanding and understanding of the environment. The term is so inclusive that it overlaps significantly with self-awareness and social awareness, because any knowledge originates from an environment, and it must be gained through the agent itself. However, given the variety of shapes and purposes of machines and models, there are a number of situations in which we need to make distinctions. Collision avoidance, responsive adaptation to dynamic changes, and accurate state estimation all represent a segment of knowledge of the environment, yet the agents will not possess knowledge about themselves. For example, an autonomous car’s awareness is seen in its continuous updating of a world model (other cars, lanes, pedestrians) and predicting future states to avoid. A more cognitive example is an AI security system being aware of the context (it “knows” that the noise it hears is the wind, not an intruder, because it integrates various sensory cues). With embodiment, sensorimotor awareness of the environment in AI can also include internal senses: a software agent might be “aware” of its CPU usage or memory limit as part of its situational context. In all cases, the hallmark is that the system maintains an up-to-date model of the environment and uses this model to guide immediate actions.
Social Awareness
The understanding of interpersonal relationships is a basis for self-formation [11]. Social awareness refers to the capacity to perceive and interpret the mental states, intentions, and social cues of others and respond effectively in a social context. Key components include theory of mind (understanding that others have independent beliefs and desires), perspective-taking (adopting others’ viewpoints), and empathy (sharing or understanding others’ emotions) [32, 33, 34]. By around age four, typical children pass false-belief tests, demonstrating theory of mind [35], whereas autistic children often fail such tasks [36]. Many animals also exhibit rudimentary social awareness, including primates [32] and birds [37]. Humans further excel at shared intentionality—collaboratively understanding and aligning with others’ goals and perspectives [38].
In artificial intelligence, social awareness can be defined as an AI system’s ability to model and respond to human mental states and social cues. The AI’s awareness of other agents (human or AI), their presence, state, and possibly their thoughts or intentions, constitutes its understanding of the surrounding society. This ranges from simple forms to advanced theory-of-mind reasoning. A chatbot noticing that a user is frustrated from their tone grasps an external understanding of human agents, while a robot deducing that its human partner is unaware of a certain fact and thus offering that information represents a higher level of societal understanding. Recent LLMs and chatbots have shown striking (if superficial) social-cognitive abilities: they can solve many theory-of-mind tasks and interpret indirect cues in conversation, sometimes approaching human-level performance [39, 40, 41] and outperforming humans on certain social reasoning benchmarks, such as choosing appropriate behavior in situational judgment scenarios [42].
@@
Table 1: Examples of Detailed Awareness Types Mapped to Core Categories
Detailed Awareness Mapped Core Category Reason
Moral/Ethical Awareness Self‑awareness + Meta‑cognition Self‑awareness: knows ethical/legal constraints;
Meta‑cognition: monitors responses for ethical risks.
@@
Spatial/Temporal Awareness Situational Awareness Focused perception, understanding and prediction of external space and time dynamics.
Emotional Awareness Social Awareness + Self‑awareness Social Awareness: perceives and responds to others’ emotions;
Self‑awareness: aware of the emotional impact of its own outputs.
Goal/Task Awareness Situational Awareness + Meta‑cognition Situational Awareness: understands task environment and progress;
Meta‑cognition: monitors effectiveness of strategies.
@@
Safety/Risk Awareness Meta‑cognition + Self‑awareness Meta‑cognition: identifies potential errors or risks;
Self‑awareness: knows its safety/compliance boundaries.
2.2 Uncovered and Overlapping sections
By definition, AI agents are able to acquire awareness about any perceived object as long as they possess the ability to receive and remember information and let us realize their knowing stage. Thus, we can trivially subdivide these categories based on common sense. For instance, goal awareness represents the knowledge of why an agent is doing something. Even relatively simple AI planners have this implicitly (the goal state is defined), but an aware AI can explicitly reason about its goals (“I need to achieve X, and I am currently pursuing sub-goal Y”). A scheduling AI might say, “I cannot schedule that meeting because I am already tasked with a higher priority deadline at that time,” showing awareness of its hierarchy of goals. Our taxonomy, as shown in Fig 1, based on the type of object and the locus of information source (internal or external), basically encompasses objects of awareness that fit the attempt, while dispatching some parts that are difficult to identify in machines (for instance, emotional awareness).
These categories are not completely independent—they often overlap, as we show in Tab 1. For example, for an AI to have social awareness, it helps if it has self-awareness (to analogize others to itself). Goal awareness and self-awareness intertwine when the AI contemplates its own goals (“I want X” entails an “I” and an objective). Nonetheless, this breakdown is useful for analyzing which aspects of awareness a given AI system has. A self-driving car has excellent sensorimotor situational awareness but no self-concept beyond coordinates on a map. A dialogue agent might have some social awareness (it tracks the user’s preferences) but zero embodiment or sensorimotor awareness. A research robot might have a self-model and thus self-awareness, but only rudimentary social understanding. The long-term vision in AI research is sometimes described as achieving integrated awareness: combining these types so that an AI can fluidly understand itself, its goals, its world, and others - a step toward more general forms of intelligence.
3 Evaluating AI Awareness in LLMs
@@
Subject Meta‑Cognition Self‑Awareness Social Awareness Situational Awareness
@@
Social robots Low Low High Low / High
@@
Whereas the scholarly discourse on machine consciousness has been anchored in ontological speculation and high‑level theoretical modeling, contemporary research on AI awareness is characterized by an epistemic commitment to empirically observable phenomenology and rigorously operationalized behavioral assays. Similar to the Turing test for testing the language intelligence of AI [3, 43], Researchers have proposed and carried out a large number of evaluation methods and work in the four aspects of self-awareness [44, 45], social awareness [46, 40, 47, 48], situational awareness [31, 49]. In this section, we specifically constrain our assessment of machine consciousness to LLMs rather than artificial intelligence more broadly for two principal reasons. First, as elaborated in Tab 2, LLMs constitute the first class of artificial agents empirically demonstrated, under controlled conditions, to exhibit all four key dimensions of consciousness. Second, to avoid conflating intrinsic model capabilities with extrinsic performance enhancements, such as retrieval modules, tool plug-ins, or multimodal interfaces, we deliberately limit our analysis to bare models. This narrower scope ensures that evaluation metrics directly reflect the endogenous mechanisms and inherent constraints of the LLM itself, rather than artifacts introduced by external augmentation, thereby yielding results more conducive to rigorous theoretical interpretation and subsequent model advancement.
3.1 Evaluation of Meta-Cogniton
Investigating the meta-cognitive abilities of large language models provides an essential perspective on the extent to which artificial systems can emulate, or even transcend, human-like reasoning and self-reflective capacities. A growing body of research demonstrates that prompting models to articulate intermediate reasoning steps—rather than directly producing answers—significantly enhances their performance on complex tasks such as multi-step mathematical problem-solving, algorithmic reasoning, and program synthesis [50, 51, 52]. This “reasoning-before-answering” paradigm, i.e., Chanin-of-Though (CoT), not only improves accuracy but has also emerged as a standard practice in the training of state-of-the-art LLMs [53, 54], suggesting its strong connection to the elicitation of metacognitive processes. Beyond performance gains, recent studies suggest that LLMs possess an emergent capacity for self-evaluation and behavioral introspection. For example, Betley et al. [55] shows that models fine-tuned exclusively on high-risk domains, such as generating insecure code or making ethically sensitive economic decisions, can independently identify and label their outputs as “dangerous,” even without explicit instructions or adversarial prompting. This spontaneous self-attribution indicates an internalized, task-specific risk-awareness characteristic of metacognitive systems. An additional line of work has highlighted the increasing sophistication of metacognitive capabilities in frontier models. In interactive or agent-like environments, models have been shown to self-reflect, identify earlier errors, and revise their responses to improve factual accuracy and task completion rates. Taking this further, research by Anthropic’s Interpretability team reveals that Claude-3.5-Haiku exhibits advanced latent-space planning during language generation [56]. Specifically, when asked to compose poetic verses, the model first anticipates suitable rhyme words at the end of a line and then generates preceding content accordingly. Rather than relying on purely token-by-token generation, it demonstrates goal-directed planning behavior, an archetypal feature of meta-cognition, where the model produces language and governs how and why confident linguistic choices are made. These results indicate that meta-cognition in LLMs is not an incidental artifact, but a structured, multi-faceted capability—manifesting in self-evaluation, error correction, risk awareness, and even latent-space planning.
@@ 3.2 Evaluation of Self-Awareness
Current AI systems, particularly LLMs, frequently self-identify as “AI assistant” and communicate using first-person pronouns. This anthropomorphic presentation has motivated extensive academic evaluations investigating deeper aspects of AI self-awareness and self-knowledge [57, 58, 59, 60, 61, 62]. To systematically assess LLMs’ awareness of their own existence and identity, Laine et al. [31] constructed the Situational Awareness Dataset (SAD), which examines LLMs’ knowledge regarding self-referential attributes, such as their model names, parameter counts, API URLs, and specific details during the training process. Their findings indicate that even the highest-performing model evaluated—Claude-3-Opus—achieved scores at approximately two-thirds of the theoretical maximum. Furthermore, these models exhibited limited capability in recognizing their own descriptive attributes. Inspired by the classical mirror test paradigm, Davidson et al. [63] further explored AI self-consistency by prompting models with self-description queries, such as providing detailed accounts of hypothetical personal experiences (e.g., attending a concert in a large stadium). Their experiments revealed significant difficulties among models in accurately identifying their own responses from multiple model-generated alternatives, highlighting a notable lack of self-consistency. Parallel research has focused on evaluating AI models’ recognition of their knowledge boundaries. Yin et al. [45] assessed models’ confidence in responding to questions that either fell beyond their trained knowledge or had no definitive answers. Results demonstrated that GPT-4 reached an accuracy rate of 75.47%, approaching human baseline performance at 84.93%. These findings suggest that although current AI systems likely do not possess human-level self-awareness, they nonetheless demonstrate preliminary levels of self-identity recognition and an awareness of their knowledge limitations.

### v2

@@ 2.1 Major Types of Awareness
Situational awareness refers to the perception, comprehension, and projection of environmental elements and their future status [96, 6, 97, 98, 99]. Endsley [100] formalized SA as “the perception of the elements in the environment within a volume of time and space, the comprehension of their meaning, and the projection of their status in the near future.” This three-level model (5b provides a thumbnail of its structure) has become the de facto definition of SA across domains: perception defines situations by tagging environmental elements semantically, comprehension integrates information, and projection supports planning and option evaluations [98]. Human situational awareness has been extensively studied using both objective and subjective measures in aviation [100, 101, 102], military [103], medical care [104, 105], and traffic circumstances [106, 107]. For objective measures, in simulated aviation battles, Endsley [100] monitored subjects’ knowledge about their location, heading direction, altitude, weapon, and information regarding their enemies, utilizing the Situational Awareness Global Assessment Technique (SAGAT) to probe operator knowledge through real-time queries during task interruptions. They integrated subjective self-reported rating scales as well for complementary reflection items. Taylor [102] developed a holistic version of the self-report instrument, Situational Awareness Rating Technique (SART), to evaluate perceptions of environmental stability, complexity, variability, etc.
EPfforts to replicate or approximate artificial situational awareness in AI systems involve enabling AI to perceive their environment, contextualize sensory data, and anticipate future events [108]. This typically involves integrating multi-sensor data into a coherent, continuously updated workplace [30]. AI-driven frameworks for situational awareness now incorporate semantic knowledge bases and real-time inference engines to track both internal system states and external environmental cues [109, 110]. For instance, an autonomous vehicle uses situational awareness to monitor nearby vehicles, interpret road conditions, and predict hazards [6, 108, 111], thereby facilitating adaptive and safe decision-making.
Given the variability of manifestations across psychology, engineering, and cognitive ergonomics [112, 99, 113], defining a strict boundary for situational awareness remains challenging yet necessary. By design, AI agents operate within predefined scenarios and possess an embedded awareness of such contexts, which often conflates aspects of self and environmental awareness. Broadly attributing behavioral changes to situational awareness risks circularity in explanation [97]. Nevertheless, capabilities such as collision avoidance, dynamic adaptation, and state estimation exemplify environment-focused situational knowledge without implying self-reflective or socially aware capacities. We delineate two concepts by confining situational knowledge to information sources that are not inherently tied to any single agent or social collective. A more cognitively rich example is an AI surveillance system that integrates audio and visual data to infer that a detected noise is caused by wind rather than a human intruder. In some cases, sensorimotor embodiment allows internal metrics, such as CPU load or memory status, to be integrated as part of an agent’s situational model. In essence, the defining characteristic of situational awareness constitutes the internal representation of the external world that enables informed decision-making, particularly in complex and dynamic operational contexts.
Decomposing awareness into metacognition, self-, social, and situational forms provides a tractable framework for evaluating and engineering intelligent systems, i.e., transforming a once vague concept into a practical research agenda.
@@
Table 1: Examples of other awareness types mapped to core categories. For brevity, we use abbreviated terms: Meta for metacognition, Self for self-awareness, Social for social awareness, and Situational for situational awareness
Other Component Reason
Moral/Ethical Self + Meta Self: knows ethical/legal constraints;
Meta: monitors responses for ethical risks.
@@
Spatial/Temporal Situational Focused perception, understanding and prediction of external space and time dynamics.
Emotional Social + Self Social: perceives and responds to others’ emotions;
Self: aware of the emotional impact of its own outputs.
Goal/Task Situational + Meta Situational: understands task environment and progress;
Meta: monitors the effectiveness of strategies.
@@
Safety/Risk Meta + Self Meta: identifies potential errors or risks;
Self: knows its safety/compliance boundaries.
@@
Subject Metacognition Self‑Awareness Social Awareness Situational Awareness
@@
Social robots No Low High Low / High
@@
2.2 Theoretical Strengths and Challenges
The adequacy of this taxonomy allows for explanations of more nuanced forms of awareness through combinations of these fundamental categories. Table 1 exemplifies that the main components adequately cover several frequently mentioned types of awareness: emotional awareness arises from perceiving one’s emotions (self-awareness) and those of others (social awareness); moral or ethical awareness involves evaluating the consequences of actions and making value judgments, thus integrating metacognition and self-awareness [114]; context awareness involves recognizing environmental spatial and temporal structures [115]. Whether some categories may overlap or not is still under debate. For instance, notwithstanding that we manually segregate “self-oriented knowledge” and “knowledge of knowing”, the intersectionality of metacognition and self-awareness depends on the rubric and paradigm of research [59]. Meanwhile, awareness studies encounter the hardship of definition vagueness, lack of unified objective indicators for evaluation, challenges posed by inconsistent interdisciplinary frameworks and objectives, and ethical concerns—we will elaborate in the following sections.
Despite being controversial, LLM dialogue systems are demonstrating a more complete awareness structure. As shown in Table 2, they exhibit a broader spectrum of cognitive capacities than robots designed for specific functions and even surpass those of some animals. They demonstrate robust mental-state reasoning in text, perform significantly better on general abilities than animals, and even exhibit advanced cognitive capacities that require profound understanding of the knowledge in their awareness pool, such as deception [116, 117, 118]. By properly regulating its strengths and weaknesses, they may have the potential to explore comprehensive awareness. In the following sections, we will explore how researchers have constructed criteria and evaluation methods to measure LLM’s capacity in “being aware of everything”.
A principled taxonomy of awareness, spanning metacognition, self-awareness, social awareness, and situational awareness, provides not only a foundation for empirical research, but also a roadmap for building more general, adaptable, and transparent AI systems. Understanding the interplay and boundaries among these dimensions is crucial both for scientific advancement and the responsible development of AI.
3 Evaluating AI Awareness in LLMs
Figure 6: Representative literature across the evaluation of major awareness dimensions
@@ 3 Evaluating AI Awareness in LLMs
Building on the preceding theory section, which defined AI awareness as a functional construct encompassing the four core types, we now turn from “what it is” to “how we measure it.” Similar to the Turing test for testing the language intelligence of AI [20, 119], researchers have proposed and carried out a large number of evaluation methodologies and studies in the four main dimensions of AI awareness, i.e., self-awareness [120, 121], social awareness [90, 118, 122, 123], situational awareness [124, 125], Figure 6 shows part of them. In this section, we specifically constrain our assessment of AI awareness to LLMs rather than artificial intelligence more broadly for two principal reasons. First, as elaborated in Table 2, LLMs constitute the first class of AI agents empirically demonstrated, under controlled conditions, to exhibit all four main dimensions of awareness to a certain level. Second, to avoid conflating intrinsic model capabilities with extrinsic performance enhancements, such as retrieval modules [126, 127], tool plug-ins [128, 129], or multimodal interfaces [130, 131], we deliberately limit our analysis to bare models, i.e., OpenAI’s o1 [132], Anthropic’s Claude-3.5-Sonnet [133], Deepseek’s R1 [134]. This narrower scope ensures that evaluation metrics directly reflect the endogenous mechanisms and inherent constraints of the LLM itself, rather than artifacts introduced by external augmentation, thereby yielding results more conducive to rigorous theoretical interpretation and subsequent model advancement.
Table 3: Summary of literature on metacognition evaluation
Authors Key Contribution Focus on Meta- -cognition Using Human Baseline Code Links
Didolkar et al. [7] Elicits GPT-4 to tag, cluster, and exploit its own math “skill” taxonomy; shows that self-selected skill exemplars boost GSM8K and MATH accuracy, demonstrating explicit metacognitive knowledge. N/A
Betley et al. [135] “Behavioral self-awareness” probes: models describe latent policies (risk-seeking, back-doors, insecure coding); touches meta-knowledge of their learned behaviors. GitHub repo
Hagendorff and Fabi [136] Latent-space Stroop-style benchmark quantifies silent “reasoning leaps” between prompt and first token—measures internal reasoning without CoT. OSF repo
Zhang et al. [137] Survey unifying Chain-of-Thought mechanisms and agent memory/perception loops; discusses meta-reasoning but is mostly a review, not an eval metric. GitHub repo
Wei et al. [138] Introduces Chain-of-Thought prompting that lets models externalise intermediate reasoning; improves tasks but is not itself metacognition evaluation. N/A
Wang et al. [139] Propose DMC: a failure-prediction + signal-detection framework that decouples metacognitive ability from task performance, yielding a model-agnostic score and showing stronger metacognition correlates with lower hallucination rates. GitHub repo
Team [140] Shows Claude-3.5-Haiku first chooses rhyme words, then fills lines—evidence of forward planning, instead of just predict next token (word). N/A
3.1 Evaluation of Metacognition
Evaluating the metacognitive abilities of LLMs provides a critical window into their capacity for introspection, self-regulation, and strategic reasoning—key ingredients of higher-order cognitive function. Following the classical three-stage framework of metacognition—(i) planning, (ii) monitoring, and (iii) evaluation (as illustrated in 4a)—recent research has begun to map how these capabilities emerge and manifest in large-scale foundation models.
  • Planning. Strategic control over generative behavior is a hallmark of advanced metacognition. While LLMs do not engage in planning through embodied trial-and-error, recent evidence suggests they can execute structured, multi-step generation pipelines internally. Anthropic’s interpretability study of Claude-3.5-Haiku [133], for example, finds that the model engages in latent planning when composing poetry: it first selects rhyming end-words, then retroactively fills in preceding lines to satisfy those constraints [140]. This mirrors human compositional planning and indicates that models may develop internal task scaffolds, even in domains that lack formal structure. Similarly, in complex reasoning tasks, models often implicitly formulate high-level response structures before surface realization, as observed in long-form summarization [141], code synthesis [142], inter alia.
  • Monitoring. Metacognitive monitoring denotes a system’s capacity to observe and assess its own cognitive operations. In LLMs this surfaces as on-the-fly self-evaluation during generation. Betley et al. [135] show that models fine-tuned on high-risk domains—e.g., insecure code or sensitive financial advice—spontaneously flag hazardous outputs, while Ji-An et al. [143] further demonstrate, via a neurofeedback paradigm, that LLMs can read out and even steer selected internal activation directions. Together, these findings suggest that models can internalise domain-specific failure patterns and respond with cautious, self-corrective framing.
  • Evaluation. Reflective reasoning—evaluating the correctness or coherence of one’s outputs—is perhaps the most studied metacognitive faculty in LLMs. Chain-of-Thought (“reasoning-before-answering,” i.e., CoT) prompting has been shown to substantially enhance performance across a wide range of reasoning tasks, from multi-step mathematics to program synthesis [138, 137, 144, 136, 7]. Consequently, CoT prompting is now baked into the training and alignment pipelines of foundation models [132, 134], underscoring its tight coupling with metacognitive processing.
Although the above work is mainly qualitative research, recently, Wang et al. [139] proposed a decoupled metacognition score that separates failure prediction from task accuracy, providing a model-agnostic gauge of self-monitoring. As shown in LABEL:tab:meta-eval-long, most studies still rely on qualitative evidence, and systematic human-baseline comparisons are lacking. Building large-scale human reference benchmarks will be crucial to understanding how architecture, scale, and training influence metacognitive capacity in future AI systems.
Table 4: Summary of literature on self-awareness evaluation
Authors Key Contribution Focus on Self- Awareness Using Human Baseline Code Links
Yin et al. [121] Assessed models’ confidence in responding to questions beyond their knowledge or without definitive answers via the SelfAware benchmark. GitHub repo
Laine et al. [124] Introduces the SAD benchmark; while targeting situational awareness in general, its Self-Knowledge subset (FACTS, INTROSPECT, SELF-RECOGNITION) partially evaluates LLM self-awareness. GitHub repo
Liu et al. [145] Think–Solve–Verify (TSV) pipeline; studies trustworthiness & introspective reasoning, incl. N/A
Cheng et al. [146] Builds model-specific Idk dataset; trains chat LLMs to refuse unknowns, mapping knowledge quadrants. GitHub repo
Tan et al. [147] ‘First-Generate-Then-Verify” framework; gauges whether a model can solve its own self-generated questions. N/A
Kapoor et al. [148] Shows fine-tuning on graded correctness yields calibrated ‘I don’t know” confidence usable in open-ended QA. GitHub repo
Chen et al. [149] Universal Self-Consistency (USC) for answer selection; improves quality but is a reasoning aid, not an SA metric. N/A
Davidson et al. [150] ‘Security-question” protocol to test self-recognition across 10 LLMs; find no robust self-ID. GitHub repo
Binder et al. [151] Shows GPT-4, GPT-4o, Llama-3 can introspectively predict their own future outputs better than other models can. HuggingFace repo
Tamoyan et al. [152] Linear-probe evidence that factual self-awareness (know / forget attributes) is encoded during generation. GitHub repo
@@ 3.2 Evaluation of Self-Awareness
Since contemporary LLMs frequently self-identify using first-person pronouns (e.g., “As an AI assistant, I…”) and already exhibit promising levels of situational and social awareness, evaluations of their self-awareness predominantly focus on deeper and subtler facets beyond basic self-referencing [151, 145, 146, 147, 148, 149, 152]. Recent assessments specifically target: (i) self-identity recognition, (ii) consistent self, and (iii) awareness of knowledge boundaries. Conceptually, these facets align with the concentric self-model, wherein self-identity recognition corresponds to the narrative self, while consistent self and knowledge-boundary awareness map onto the minimal self. The innermost layer, no-self (absence of self-identification), is typically not evaluated, as modern LLMs inherently surpass this baseline through their self-referential dialogue.
  • Self-Identity Recognition. The Situational Awareness Dataset (SAD [^6] A benchmark designed to assess various dimensions of model awareness, including but not limited to self-awareness. It includes subsets targeting self-knowledge (e.g., model name, size, training details) as well as broader situational understanding. It should not be confused with the models under evaluation.) [124] examines whether models know details about themselves, such as their name, parameter count, API endpoints, and training specifics. Even top-performing models, such as Claude-3-Opus [153], achieve only about two-thirds of the theoretical maximum and show limited capability in detailed self-description.
  • Consistent Self. Inspired by the mirror test, Davidson et al. [150] prompt models to distinguish their own past responses from distractors. Models often struggle to accurately identify their previous outputs, particularly when responding to prompts involving vivid yet hypothetical experiences, indicating limited internal coherence.
  • Knowledge-Boundary Awareness. Confidence calibration studies [121, 154] show that GPT-4 identifies whether it knows the answer to ambiguous or unanswerable questions with 75.5% accuracy—approaching but still below the human baseline of 84.9%, i.e., LLMs show a relatively clear knowledge-boundary.
Overall, according to LABEL:tab:self-eval-long, contemporary LLMs demonstrate initial capabilities in narrative and minimal self-awareness, although they remain distant from human-level self-reflection and robust coherence across diverse contexts. Future work should further explore neglected aspects of LLMs’ self-awareness, including minimal self-autonomy, the stability of self-descriptions across varying contexts, and sustained cross-turn coherence, to build a more comprehensive understanding of this topic.
Table 5: Summary of literature on social awareness evaluation
Authors Key Contribution Focus on Social Awareness Using Human Baseline Code Links
[…]
Jiang et al. [155] Builds Commonsense Norm Bank (1.7M moral judgements) and trains Delphi, which hits 92.8% agreement with human crowd labels—beating GPT-3 (60%) and GPT-4 (79%)—thereby benchmarking LLM moral-norm awareness. N/A
Qiu et al. [156] Created cross-cultural norm benchmark; finds GPT-4 violates 12 % of norms vs 4% human, GPT-3 violates 28%. GitHub repo
Voria et al. [157] Presents first SE-oriented framework mapping developer-side ethics vs runtime collaboration; outlines future evaluation axes. N/A
Li et al. [158] Proposed five-factor awareness taxonomy; among 13 LLMs, social-awareness tops at 78% (GPT-4) whereas capability-awareness stays at 40%. GitHub repo
Zhuge et al. [159] Assembles up to 129 agents in a Natural-Language Society-of-Mind; VQA accuracy rises to 67% vs 60% best single model, showcasing emergent multi-agent social reasoning across multimodal tasks. GitHub repo
Choi et al. [160] Released 4k-scenario SOCKET dataset; shows GPT-4 matches crowd sentiment/offense judgements (85%) but trails on trust, GPT-3 lags by 20 pp overall. GitHub repo
Xu et al. [161] Introduced six interactive tasks; Chain-of-Thought lifts GPT-4 success to 63% yet 30% failures persist under uncertainty, GPT-3 <<<25%. HuggingFace repo
Gandhi et al. [162] Built higher-order ToM benchmark; reveals GPT-4 accuracy crashes below 10% on second-order beliefs, GPT-3 at chance. GitHub repo
[…]
Li et al. [164] Introduced role-playing “AI Society” (100 k dialogues); GPT-4 collaborative success ↑20 pp over single-role chats, indicating improved cooperative social reasoning. GitHub repo
Park et al. [165] Simulated a 25-agent “small-town” sandbox; human raters judged 81% of agent actions socially plausible, showing memory + reflection + planning yields emergent social behaviour. N/A
Rao et al. [166] Launched 11-language norm dataset; uncovers 25 pp drop for GPT-4 on Global-South norms, few-shot tuning recovers 15 pp. GitHub repo